### v1

# Analysis of optical pattern formation on glass

Abstract
We investigate the formation of particular patterns when light passes through glass. Experimentally, we use various glass plates, registering each reflected and transmitted outcome. Thus, we find the condition to form a common pattern, namely, randomly scratched plates produce the halos. We engineer other optical patterns by means of the specular holography method. Using the Defocusing Microscopy technique, we study the geometric properties of the glass, establishing the physical and mathematical model for a simulation. Our computational results agree well with the experimental ones, leading us to conclude that the pattern formation is governed by specular reflection on various points along the scratches. Our findings give an initial explanation for the pattern formation on glass, in particular for the halo design.
Keywords: Glass halo, optical distortions, optical patterns, geometric optics, International Physicists’ Tournament
1 Introduction
In optics, the interaction of light with rough surfaces is fundamental to understanding everyday effects, considering that objects are not perfectly smooth. For instance, the light scattering phenomenon involves the redirection of light by oscillating charges and it approaches the limit of geometric optics for large particles in comparison to the wavelength of light [1]. The scattered light intensity and profile were predicted by various complex methods [2, 3, 4], but results reported in the literature indicate that the theory is not yet complete [5, 6]. Interestingly, the surface roughness can be used to create holograms. Experiments with plastic paper and a compass revealed a technique, called specular holography, for creating holograms by hand [7, 8]. The arched scratches on the surface act as a curved mirror and form the tridimensional (3D) image. A computational method successfully predicted the bright spots that compose the hologram by applying geometric optics and specular reflection [9], which greatly simplify the light-surface interaction in this scope.
Previous works mainly focused on developing broad theories for the interaction between light and surface roughness. However, the unguided formation of certain patterns on the surface of reflective materials, mainly glass, remains to be described. In this paper, we start from the problem “Glass halo”, proposed for the International Physicists’ Tournament (IPT) 2023 [10]. The problem statement goes as: “Glittering circles can be seen when light from a source with small angular size passes through a glass. On closer examination they appear to be composed of small scratches and structural inhomogeneities. In some cases, specific rays can be seen, diverging from the light source. Under which conditions such circle halos and lines can be seen? Investigate their geometrical properties and what shapes you can engineer.” Accordingly, we started by experimenting with microscope glass plates and registering the pattern outcome. A geometric investigation laid the basis for a simulation that agrees with the geometric optics simplification proposed by [9]. As we will show, the resulting optical design allows us to specify the roughness pattern in various materials. This identification is an initial progress when dealing with everyday objects that do not always appear rough to the human eye; for instance, the patterns discussed here can be seen in glass windows, cellphone screens and car surfaces. From the viewpoint of applications, the understanding of pattern formation associated with roughness improves studies of surface polishing.
The remaining sections are organized as follows. In Sec. 2, we present all methods: experimental procedures with glass plates, simulation, and experimental process with the specular holography technique. In Sec. 3, we present and discuss the experimental results, comparing them to the simulation when suitable, and discuss the points for additional studies. Finally, our concluding remarks are stated in section 4.
2 Methods
2.1 Experiments
We started our investigation with experiments varying two factors: (a) the glass surface and (b) the distance between the glass plate and the camera-light source combination: see Fig. 1a for our experimental setup. We used a polished as well as mechanically scratched glass plates. The materials and their specifications are listed in Table 1. The cellphone flashlight is composed of light emitting diodes (LEDs) and its intensity was constant during all the experiments carried out for this paper.
Item Characterization
Plate 1 Polished glass plate
Plate 2 Horizontally scratched plate with a 150 grit sandpaper
Plate 3 Vertically scratched plate with a 150 grit sandpaper
Plate 4 Randomly scratched plate with a 80 grit sandpaper
Plate 5 Randomly scratched plate with a 150 grit sandpaper
Plate 6 Randomly scratched plate with a 280 grit sandpaper
Flashlight Cellphone flashlight with constant intensity
Measuring tape Milimiter tape attached to the table
Cellphone camera 12MP Dual camera with digital zoom up to 5x
Table 1: Materials used in the experiments with glass. All plates have the same dimensions: $a_{1}=(7.6\pm 0.1)$cm, $a_{2}=(2.5\pm 0.1)$cm, and $a_{3}=(0.10\pm 0.05)$cm, where $a_{1}$ is the height, $a_{2}$ is the width, and $a_{3}$ is the thickness.
During the first part, we took photos of the reflected (observer 1) and transmitted (observer 2) patterns for each glass plate (see again Fig. 1). In all these cases, the distance between the glass surface and the flashlight was $d_{f}=(0.55\pm 0.01)$m. Similarly, the distance between the surface and both cameras (observers) was always $d_{c}=(0.55\pm 0.01)$m.
In the second round of experiments, we varied the distances $d_{f}$ and $d_{c}$ for plates 5 and 6, taking photos of the reflected and transmitted patterns. Starting at $d_{f}=d_{c}=(0.55\pm 0.01)$m, we reduced ten centimeters at a time until $d_{f}=d_{c}=(0.25\pm 0.01)$m.
Figure 1: (a) Sketch of the experimental setup. The distances from the glass plate to the flashlight $d_{f}$ and to the camera $d_{c}$ at either positions (viewer 1 or 2) were equal. (b) Glass plates described in Table 1. Since the light source is not perpendicular to the scratches, we do not see the formation of optical patterns.
Subsequently, we studied the geometrical properties of the glass plates in order to characterize the scratches properties. In this context, the defocusing microscopy technique [11] is used to reconstruct the thickness and curvature profiles of phase objects, such as red blood cells [12] and waveguides [13]. We used it to reconstruct the profile of the scratches on selected regions of plates 4, 5, and 6. For cases that only involve refraction and reflection, the use of geometric optics is more than sufficient. Thus, we checked whether the scratches have dimensions greater than the wavelength of light.
2.2 Simulation
Following Ref. [9], our simulation recreates the reflection of light on scratched glass based on two main assumptions:
    in our experiments, the reflected and transmitted outcomes are equal;
    the physical regime is that of geometric optics.
The first one is correct for a thin plate or when the observer is far from the surface, which are two conditions satisfied in our configuration - see [9] for a detailed discussion. Although the transmission and reflection coefficients are distinct and the intensity will differ, the pattern for both observers is the same. The second premise is a simplification that we will argue to be valid for this phenomenon based on (i) the good agreement between our experiments and simulations and (ii) the investigation of scratch size using defocusing microscopy. Consequently, we only analyze the reflected pattern and use the law of reflection, given by
$\hat{E}(t)\cdot\hat{q}(t)=\hat{R}(t)\cdot\hat{q}(t),$ (1)
to plot the light rays. In equation (1), $\hat{E}(t)$ is the unit vector in the direction of the incoming ray, that is, corresponding to the vector that goes from the light source to a given point on the scratch, while $\hat{R}(t)$ defines the unit vector in the direction of the outgoing ray, i.e., the direction of the vector that goes from the point on the scratch to the camera (see Fig. 2). The unit vector $\hat{q}(t)$ defines the tangent line to the scratch curve.
Figure 2: Simulation outline. The incoming ray reaches the scratch and is reflected towards the camera. The reflection is seen (plotted) if it obeys the tolerance criterion.
The scratches on everyday objects are random unless it undergoes an ordered process. Also, we assume that the scratches behave like a gap in the glass with semicircular cross section and are locally linear. This shape produces the following feature: the reflected ray can exit at any angle in the plane passing through the section, since it depends on the point that the light hits the semicircle. This implies that a scratch can reflect the ray at any angle in a plane passing through the point, which is a reasonable assumption, considering that we do not know the profile of the scrape. Lastly, the light source is considered point-like.
Equation (1) can be written as
$|\hat{E}(t)||\hat{q}(t)|cos(\theta_{1})=|\hat{R}(t)||\hat{q}(t)|cos(\theta_{2}),$ (2)
where $\theta_{1}$ is the angle between $\hat{E}(t)$ and $\hat{q}(t)$ and, similarly, $\theta_{2}$ is the angle between $\hat{R}(t)$ and $\hat{q}(t)$. For the camera or our eyes to capture the reflected beam the following relation must be satisfied, within a narrow tolerance,
$cos(\theta_{1})-cos(\theta_{2})\leq TOL.$ (3)
In our simulation, written in Python, first we generate the scratches (randomly or ordered) and then test the law of reflection for every point along the scratches. Only points that obey the cosine difference of Eq. (3) for a given tolerance ($TOL=0.0125$) are visible in the plotted results.
We set the simulation parameters to match our experimental arrangement. Initially, the light source and the camera are located at $\vec{r}=(0.55\mbox{ m})\hat{z}$ (in the coordinate system defined in Fig. 2, for which $\hat{x}$ and $\hat{y}$ are in the plate plane), corresponding to the distances $d_{f}=d_{c}=(0.55\pm 0.01)\,$m. At this configuration, we simulated the outcome for the scratched plates, varying the number of scratches on the plate for the random case. Lastly, with a defined number of random scratches, we moved the camera and light positions in the simulation ten centimeters at a time until reaching $d_{f}=d_{c}=(0.25\pm 0.01)\,$m as in the experiments.
Even though our simulation was based on a holography paper [9], it is still valid, considering that the study is based on simple reflection and refraction laws, and we generate linear scratches to match the experiments, deferring from the reference, which uses scratches curved in a specific manner to form holograms - see Sec. 2.3.
2.3 Specular Holography
Besides the pattern formation of each glass plate, we relied on the specular holography technique [7, 8] to engineer new shapes. Although this technique can be applied to glass, we used plastic surfaces (CD cases; usually made of polystyrene [14]) since it was easier to manipulate and scratch. The materials for this experimental part are recorded in Table 2.
Item Purpose
Paper Used to draw the guiding blueprint
School compass Used to scrape the plastic
Jewel CD cases Scratching surface
Cellphone camera 12MP Dual camera with digital zoom up to 5x
Table 2: Materials used while applying the specular holography technique.
We start by drawing the desired shape on the paper that acts as the blueprint. Then, the fixed portion of the compass is placed on the blueprint, and we rotate the compass gently scratching the plastic surface with its metal leg (see Fig. 3). We apply this procedure to various points along the blueprint, always marking the plastic in a circular motion. The compass aperture angle was constant at $\phi=(35\pm 1)$°.
Figure 3: Schematic figure of the drawing process. One leg of the compass stays on the blueprint, following the points along the desired shape.
We visualize these engineered shapes by placing the flashlight and the camera at, respectively, $d_{f}=(0.94\pm 0.01)\,$m and $d_{c}=(1.02\pm 0.03)\,$m above the plastic surface. To see the hologram movement, we slightly move the camera along a horizontal plane, keeping it at a distance $d_{c}$ above the surface.
3 Results and discussions
3.1 Observed patterns
From our experiments with the various glass plates, we determine the conditions to form the main patterns, namely, (1) randomly scratched surfaces produce the halos and (2) bright spots tend to generate diverging rays, which are optical distortions.
First, the randomly scratched glass plates generate the glass halo, as shown in Fig. 4 for plates 4, 5, and 6 described in Table 1. The halos are visible for all the randomly scratched plates regardless of the sandpaper used, and, as expected, the halo lines appear more fine as we increase the sandpaper grit. We also observed the formation of this halo in other materials with randomly scraped surfaces, such as glass windows and cellphone screens. This type of surface reflects the light from all directions, and a circular design is anticipated.
Figure 4: Halo formation for plates scratched with (a) a coarse grit, (b) with a medium grit and (c) with a fine grit.
The second main pattern is the formation of diverging lines, as seen in Fig. 5. They are an optical distortion in the imaging process and it can occur to any of the plates of Table 1, since it is related to the way the camera lenses capture the strong light source. Indeed, in our experiments, we noticed that the lines are not visible to the bare eyes and that they rotate as the camera rotates (but not when the flashlight or the glass plate rotates).
Figure 5: Distortion lines seen (a) through a polished glass (plate 1), (b) reflected on a scratched plate (plate 6) and (c) directly from the flashlight. The green dot also indicates that it is an imaging distortion.
Lastly, when analyzing the reflection and refraction from polished plates, we noticed a saturated light circle surrounded by a blurred pattern that covered the entire glass surface - see now Fig. 5a. This blur is visible to the eyes and differs from the halos because it is not composed of scratches.
3.2 Scratch reconstruction
If we consider a scratched plate on the xy plane and look at it from the z axis, the width of the scratch on this plane determine the size of the obstacle/slit. Therefore, the scratch’s width is the relevant parameter to compare with the wavelength of light.
Using the defocusing microscopy method, we estimated the mean width $a$ of the scratches after reconstructing their surface and lateral profile, see Fig. 6. From the results presented in Table 3, we observe that the width tends to be larger than the wavelength range of the cellphone flashlight $\lambda_{\,light\,}=(0.400-0.750)\mu$m [15]. Also, in our experiments, we do not observe interference or diffraction patterns. These factors indicate that the geometric optics hypothesis, assumed in our simulation, is reasonable.
Plate Estimated size
Plate 4 $a=(15\pm 2)\mu$m
Plate 5 $a=(15\pm 2)\mu$m
Plate 6 $a=(16\pm 6)\mu$m
Table 3: Mean width of the reconstructed scratches for the randomly scratched plates. The uncertainties are the standard deviation from seven measurements.
Figure 6: (a) Reconstruction of plate 5. We obtained the mean width from the marked scratches. (b) Three dimensional plot of a selected reconstructed area. (c) Lateral profile of a selected scratch.
3.3 Simulation
As detailed in Sec. 2.2, the final images created in our simulation are the patterns generated when light is shone over the scratched plates [given by points that obey Eq. (3)]. Thus, our simulation provided images of the pattern designs predicted using our theoretical model, and we compared them to our experimental results.
Starting with the horizontally and vertically scratched glass (plates 2 and 3), in Fig. 7 we show a comparison between the experimental and the theoretical results. In the results, we notice that a horizontally scraped plate highlights a vertical light pattern, and vertically scratched plates produce a horizontal highlight. We see a close similarity between the predicted and the experimental patterns.
Figure 7: Comparison between (a) experimental and (b) computational results for a horizontally scratched plate, showing a vertical highlight. Similar analysis comparing (c) experimental and (d) computational patterns for a vertically scratched plate, showing a horizontal line. Both simulation trials we conducted with $d_{f}=d_{c}=0.55\,$m and a fixed number of scratches $N=400$.
Proceeding to the case of randomly scratched plates, we first compare the simulated and experimental results in Fig. 8, confirming that our model describes well our results.
The consistency between the results indicate that the simplifications of our model are valid, and we used our simulation to study a characteristic of the phenomenon, i.e., the halo density. We did it by altering two factors: the number of scrapes and the distance to the light-camera combination.
As we progressively increase the number of scratches on the surface, the halo density also increases, as we can observe in Fig. 9. It happens because for a larger number of scratches, more points obey the law of reflection for the small tolerance imposed, and the halo will appear fuller.
Similarly, for a fixed number of scratches, we see a fuller circle of halos when the plate is relatively far from the light-camera group - see now Fig. 10. As we move the surface closer to the light and the camera, fewer light rays contact the scrapes and are reflected in the desired angle of the observer.
Figure 8: Comparison between (a) experimental and (b) simulation results for the halo at $d_{f}=d_{c}=(0.55\pm 0.01)\,$m. In the simulation, the number of scratches was $N=3000$.
Figure 9: Halo result as we increase the number of scratches from (a) $N=1500$ to (b) $N=2000$ and, finally, to (c) $N=3000$. Trials conducted with a fixed distance of $d_{f}=d_{c}=0.55\,$m.
Figure 10: Halo formation for (a) $d_{f}=d_{c}=0.35\,$m, (b) $d_{f}=d_{c}=0.45\,$m and (c) $d_{f}=d_{c}=0.55\,$m. In all these trials the number of scratches was constant ($N=2000$).
3.4 Engineered patterns
We were able to engineer the two desired patterns, a square and a circle, using the specular holography method (see Fig. 11); the video showing the complete process for hologram creation is available as supplementary material.
We notice that the glass halo phenomenon does not form a 3D image. In the literature, the scratch hologram formation seems to demand scratches in a specific arch form, which act as curved spherical mirrors, producing either a real or virtual image [7, 16], while the halo was formed with random locally linear scratches.
Figure 11: Two patterns created by scratching the CD cases: (a) a square hologram and (b) a circle hologram.
3.5 Further studies
One way to improve this study is by verifying the equivalency between reflected and transmitted patterns. We used this fact to analyze only the reflected one in the simulation, but to certify it in the experiments one would need to capture the halo at the same position and angle on both sides of the surface.
Another difficulty is faced while observing the effects of distance on the halo density. The point of light incidence on the glass needs to be accurate every time we decrease the distance, ensuring that it is the only varying factor. While we tried to maintain this caution in the experiments, one would need a more precise experimental setup to ensure the results and compare the densities.
Lastly, an advancement can occur by obtaining a satisfactory result from the specular holography method on glass surfaces, rather than on a plastic material. The difficulty relies on scratching the glass in a controlled circular manner. To do that, one would need the appropriate pointer, e.g., a silicon carbide tool. An attempt to create a hologram on glass using laser engraving, as well as its limitations, was reported by Augier [17].
4 Conclusions
We investigated the formation of optical patterns on glass. We observed experimentally that randomly scratched glass plates produce the glass halo phenomenon, highlighting the structural inhomogeneities. The second pattern we studied was the formation of diverging lines, which comes from an optical distortion of the imaging process; indeed we observed that they rotate as we rotate the camera. We assumed that the phenomenon of light passing through a glass that has been mechanically scratched can be acceptably described using ray optics; indeed a simulation based on this assumption well reproduced our experiments. According to our results, the halo formation on everyday reflective objects, such as glass, is governed by specular reflection on various points along the scratches. We used our simulation to study the effect of two factors on the halo density, namely, (1) the number of scratches and (2) the distance between the surface and the camera-light source. As we progressively increase the number of scratches on the surface, the halo density also increases; similarly, we see a fuller circle of halos when the plate is relatively far from the light-camera position.
Acknowledgements
The authors would like to thank Ariel Guimarães and Maria Vargas for their contribution, especially in the first stages of research. We also thank João Pedro Cunha for his help during experiments, and Anna Luisa Lemos for producing the supplementary video. MCOA and UA acknowledge financial support from CNPq, CAPES, and FAPEMIG.